OpenStudy (anonymous):

7. Solve. 8 - 3x - x = 12

OpenStudy (anonymous):

8 - 3x - x = 12 8-4x=12 -4x=1.5 x=-0.375

jhonyy9 (jhonyy9):

8-3x-x=12 -4x=4 x=4/(-4) = -2 x=-1